# Dirac’s variational approach to semiclassical Kramers problem in Smoluchowski limit

### v1

@@ Abstract
Kramers escape from a metastable state in the presence of both thermal and quantum fluctuations under strong damping is treated as a thermally activated process in a quantum modified semiclassical potential. Dirac’s time-dependent variational method together with the Jackiw-Kerman function is employed to derive the semiclassical potential.
@@ I Introduction
The problem of escape from metastable states is ubiquitous in physics, chemistry, and biology. A particle can escape from a metastable state owing to thermal fluctuations, or quantum tunneling, or both effects.
@@
The work of Kramers K represents a classic analysis of such problems as a thermally activated processes. He treated the escape process as a Brownian motion of the particle in the metastable state driven by thermal noise. The dynamics was given in terms of some probability density function that satisfies an equation of the Fokker-Planck type, called the Kramers equation G ; R .
@@
In recent years there has been much interest in the study of quantum Kramers problem, which considers both the thermal and quantum fluctuations. This is mainly motivated by the work of Caldeira and Leggett M1 . They studied quantum tunneling in the presence of dissipation at zero temperature. Their motivation was to understand quantum mechanics at the macroscopic level as stimulated by the problem of Schrödinger cat. The method of path-integral was found useful in their work, and has since become the main tool in subsequent works by others. Extension to finite temperatures were given in M2 ; M3 ; M4 . Candidates for such macroscopic quantum phenomena include, for example, Josephson junction and SQUID circuit, single-domain magnet, and liquid Helium-3 and -4 mixture M5 . For reviews of the earlier works, see e.g., QK1 ; QK2 . Soon after, different approaches have been proposed to study the problem, e.g., by Wigner phase space W1 ; W2 , and by quantum Smoluchowski equations A1 ; A2 ; A3 ; S1 ; S2 .
@@
For quantum Kramers problem, one is basically interested in the quantum correction to the thermally activated escape rate over a potential barrier. In this case, a semiclassical approach is usually adopted. Most of the works appear so far make use of the complicated machinery of path integral. This method, while elegant, unfortunately involves continuation into imaginary time in its formalism. As such it is not easy to have an intuitive understanding of the physical process involved. Furthermore, unlike the classical theory of Kramers, which considers the evolution of probability distribution function of the Brownian motion of a particle under a force field, the path integral method deals mainly with the evaluation of quantum partition function of the system interacting with a bath of harmonic oscillators having a certain distribution of frequencies. Different distributions represent different damping characteristics. Thus the result depends very much on the frequency spectrum of the oscillators assumed.
@@
In our view, a simple and more intuitive semiclassical approach has been the time-dependent variational method initiated by Dirac D1 ; D2 . This method gives the effective semiclassical Hamiltonian very directly. Previously we have employed this method to study the phases of the ground states of the quantum Frankel-Kontorova model H1 , and the quantum chaotic behavior of the Hénon-Heiles model H2 . In these work we found it very useful to take the Jackiw-Kerman function JK ; C as the trial wavefuncrtion for variational method.
@@ II Classical Kramers escape rate
From now on we shall consider the escape problem of a metastable state from a potential $V(x)$ having a minimum $V(x_{0})=0$ at the origin $x_{0}=0$ and a maximum with barrier height $V(x_{b})>0$ at $x_{b}>0$, together with the boundary conditions $V(x)\to\infty,V_{\infty}<0$ as $x\to-\infty,\infty$, respectively. The particle, once escape, will move towards $x\to+\infty$. One such potential is the cubic potential, $V(x)=Ax^{2}-Bx^{3}(A,B>0)$. We consider the situation where $V(x_{b})\gg k_{B}T\gg\hbar\omega_{0}$, here $\omega_{0}\equiv\sqrt{V^{\prime\prime}(x_{0})/M}$ is the frequency of small oscillation around the minimum $V(x_{0})$.
@@ III Dirac’s time-dependent variational approach
$V_{sc}(x)=V(x)+\hbar\left[\frac{1}{8MG}++\frac{2}{M}G\Pi^{2}+\frac{1}{2}GV^{\prime\prime}\right].$
@@
$\langle\Psi|Q|\Psi\rangle=x,\quad\quad\langle\Psi|P|\Psi\rangle=p,$ (9)
@@
$\langle\Psi|(P-p)^{2}|\Psi\rangle=4\hbar\,G\Pi^{2}+\frac{\hbar}{4G}.$ (11)
@@ IV Effective potential and semi-classical Kramers escape rate
${\hat{V}}(x)=V(x)+\hbar\left[\frac{1}{8MG}++\frac{1}{2}GV^{\prime\prime}\right],$ (15)
@@
$r_{sc}\approx r_{c}e^{\frac{\hbar\omega_{0}}{4k_{B}T}\left(1+\frac{\omega_{b}^{2}}{\omega_{0}^{2}}\right)}.$ (18)
@@ V Discussion
The result (17) is consistent in form with those from the quantum Smoluchowski equations considered in QK1 ; A3 , except the quantum fluctuation in position $\hbar G$ is replaced by the symbol $\lambda$ in these two references. However, the viewpoints adopted in the derivations of the results are different.
@@
$D_{q}=D\left[1-\frac{\lambda V^{\prime\prime}(x)}{k_{B}T}\right]^{-1},$
where $\lambda=\langle x^{2}\rangle-\langle x^{2}\rangle_{cl}$ is the quantum fluctuation in position. The values of $\lambda$ is substantiated by path-integral computations for different damping cases and temperature regimes.
In this respect we note that an earlier version of the quantum Smoluchowski equation proposed in A1 ; A2 ; S1 ; S2 , quantum corrections appeared in both the drift potential and the diffusion term. As a result, the exponent in the quantum enhancement factor in (17) is doubled. This has been noted in W2 ; QK1 ; A3 .
VI Acknowledgments

### v2

@@ Abstract
Kramers escape from a metastable state in the presence of both thermal and quantum fluctuations under strong damping is treated as a thermally activated process in a quantum modified semiclassical potential. Dirac’s time-dependent variational method together with the Jackiw-Kerman function is employed to derive the semiclassical potential. Quantum correction is incorporated in the drift potential, and is determined by quasi-stationary conditions and minimal uncertainty relation. The semiclassical rate obtained here is consistent in form with those from the quantum Smoluchowski equations deduced heuristically by modifying the diffusion coefficient using the path-integral method. Unlike approaches using the path-integral, which involves continuation into imaginary time, the approach here is simpler and more easily understood in terms of classical picture.
@@ I Introduction
The problem of escape from metastable states is ubiquitous in physics, chemistry, and biology. A metastable state is a temporary stable non-equilibrium state of a system with a considerable long lifetime before the particle transits to a lower energy state. Common examples are the alpha-decay of radioactive atoms, supersaturated vapor, transition states in chemical reactions, etc. System in such a state can be considered as being trapped in a local minimum of the energy potential of the system. Escape from a metastable state can be triggered by thermal fluctuations, or quantum tunneling, or both effects.
@@
The work of Kramers K represents a classic analysis of such problems as thermally activated processes. He treated the escape process as a Brownian motion of the particle in the metastable state driven by thermal noise. The dynamics was given in terms of some probability density function that satisfies an equation of the Fokker-Planck type, called the Kramers equation G ; R .
@@
In recent years there has been much interest in the study of quantum Kramers problem, which considers both the thermal and quantum fluctuations. This is mainly motivated by the work of Caldeira and Leggett M1 . They studied quantum tunneling in the presence of dissipation at zero temperature. Their motivation was to understand quantum mechanics at the macroscopic level as stimulated by the problem of Schrödinger cat. The method of path-integral was found useful in their work, and has since become the main tool in subsequent works by others. Extension to finite temperatures were given in M2 ; M3 ; M4 . Candidates for such macroscopic quantum phenomena include, for example, Josephson junction and SQUID circuit, single-domain magnet, liquid Helium-3 and -4 mixture M5 , and certain phenomena in relativistic physics M6 . For reviews of the earlier works, see e.g., QK1 ; QK2 . Soon after, different approaches have been proposed to study the problem, e.g., by Wigner phase space W1 ; W2 , and by quantum Smoluchowski equations A1 ; A2 ; A3 ; S1 ; S2 .
@@
For quantum Kramers problem, one is basically interested in the quantum correction to the thermally activated escape rate over a potential barrier. In this case, a semiclassical approach is usually adopted. Most of the works appear so far make use of the complicated machinery of path integral. This method is elegant and powerful in dealing with open systems with noise and dissipation. However, it involves continuation into imaginary time, which in our opinion does not provide a direct and intuitive understanding of the physical process involved. Furthermore, unlike the classical theory of Kramers, which considers the evolution of probability distribution function of the Brownian motion of a particle under a force field, the path integral method deals mainly with the evaluation of quantum partition function of the system interacting with a bath of harmonic oscillators having a certain distribution of frequencies. Different distributions represent different damping characteristics. Thus the result depends very much on the frequency spectrum of the oscillators assumed.
@@
In our view, a simple and more intuitive semiclassical approach has been the time-dependent variational method initiated by Dirac D1 ; D2 . This method gives the effective semiclassical Hamiltonian very directly. Previously we have employed this method to study the phases of the ground states of the quantum Frankel-Kontorova model H1 , and the quantum chaotic behavior of the Hénon-Heiles model H2 (a different chaotic system has been considered by this method in E ). In these work we found it very useful to take the Jackiw-Kerman function JK ; C as the trial wavefuncrtion for variational method.
@@ II Classical Kramers escape rate
For Kramers escape problem, it is typical to consider a potential with two minima separated by a barrier, such as the asymmetric/symmetric double-well potential, or a completely asymmetric potential with a minimum and a maximum beyond which the potential is non-confining. Here, for definiteness, we shall consider a potential $V(x)$ having a minimum $V(x_{0})=0$ at the origin $x_{0}=0$, a maximum with barrier height $V(x_{b})>0$ at $x_{b}>0$, and a local minimum at $x_{m}>x_{b}$ (such as the double-well), or $V(x)\to V_{\infty}<0$ as $x\to\infty$. Without loss of generality, we assume the boundary condition $V(x)\to\infty$ as $x\to-\infty$, i.e., a reflecting boundary at $x\to-\infty$. The particle, once escape from the minimum at $x=0$ , may stay in the minimum at $x_{m}$ (in this case the particle may escape back to $x_{0}$ again), or will move towards $x\to+\infty$ without returning. The double-well is described by a quartic function, such as $V(x)=Ax^{2}(x-B)^{2}(A,B>0$). For the non-confining potential a commonly employed one is the cubic potential, $V(x)=Ax^{2}-Bx^{3}(A,B>0)$. We consider the situation where $V(x_{b})\gg k_{B}T\gg\hbar\omega_{0}$, here $\omega_{0}\equiv\sqrt{V^{\prime\prime}(x_{0})/M}$ is the frequency of small oscillation around the minimum $V(x_{0})$.
@@ III Dirac’s time-dependent variational approach
$V_{sc}(x)=V(x)+\hbar\left[\frac{1}{8MG}+\frac{2}{M}G\Pi^{2}+\frac{1}{2}GV^{\prime\prime}\right].$
@@
$\langle\Psi|Q|\Psi\rangle=x,\quad\quad\langle\Psi|P|\Psi\rangle=p,$
@@
$\langle\Psi|(P-p)^{2}|\Psi\rangle=4\hbar\,G\Pi^{2}+\frac{\hbar}{4G}.$
@@ IV Effective potential and semi-classical Kramers escape rate
${\hat{V}}(x)=V(x)+\hbar\left[\frac{1}{8MG}+\frac{1}{2}GV^{\prime\prime}\right],$ (13)
@@
$r_{sc}\approx r_{c}\exp\left\{\frac{\hbar\beta\left(\omega_{0}^{2}+\omega_{b}^{2}\right)}{4\omega_{0}}\right\}.$ (16)
@@ V Discussion
The result (15) is consistent in form with those from the quantum Smoluchowski equations considered in QK1 ; A3 , except the quantum fluctuation $\hbar G$ is replaced by the symbol $\lambda$ in these two references. However, the viewpoints adopted in the derivations of the results are different.
@@
$D_{q}=D\left[1-\frac{\lambda V^{\prime\prime}(x)}{k_{B}T}\right]^{-1},$ (17)
where $\lambda=\langle x^{2}\rangle-\langle x^{2}\rangle_{cl}$ is the quantum fluctuation in position. The values of $\lambda$ is substantiated by path-integral computations for different damping cases and temperature regimes. For instance, in the strong damping limit considered in this work, the escape rate by this approach is determined to be QK1
$r_{q}\approx r_{c}\exp\left\{\frac{\hbar\beta\left(\omega_{0}^{2}+\omega_{b}^{2}\right)}{2\pi\gamma}\left[\psi\left(1+\frac{\hbar\beta\gamma}{2\pi}\right)-\psi(1)\right]\right\},$ (18)
where $\psi(z)$ is the digamma function. This is similar to (16), except in the exponent the factor $\omega_{0}$ is replaced by $\gamma$ and a numerical factor in terms of the digamma function. The difference is of course due to the different assumptions employed in the two approaches. Computing quantum corrections by the path integral method has the advantage that it could cover different damping regimes and for non-Markovian noise as well, but the results depend very much on the assumed frequency distribution of the bath of oscillators that coupled to the Brownian particle, and continuation into imaginary time is used, which is physically less intuitive. Also, as mentioned before, the form of the quantum modified drift coefficient in this quantum Smoluchowski equation is assumed heuristically. In contrast, the semiclassical approach adopted here is more easily understood in terms of classical picture, and the quantum corrected drift potential (not the diffusion coefficient), is derived by a simple variational principle. In this respect, we note that in an earlier version of the quantum Smoluchowski equation proposed in A1 ; A2 ; S1 ; S2 , quantum corrections appeared in both the drift potential and the diffusion term. As a result, the exponent in the quantum enhancement factor in (15) is doubled. This has been noted in W2 ; QK1 ; A3 .
In this work we have considered the stochastic force on the Brownian particle to be the Gaussian noise. It should be of interest to extend the work to cases with colored noise.
Acknowledgments